### v1

# On the Relationship between Space-Time Accessibility and Leisure Activity Participation

Abstract
Understanding how accessibility shapes participation in leisure activities is central to promoting inclusive and vibrant urban life. Conventional accessibility measures often focus on potential access from fixed home locations, overlooking the constraints and opportunities embedded in daily routines. In this study, we introduce a space–time accessibility (SPA) metric rooted in the capability approach, capturing feasible leisure opportunities between home and work given a certain time budget, individual transport modes, and urban infrastructure. Using high-resolution GPS data from 2,415 residents in the Paris region, we assess how SPA influences total travel time and leisure participation, measured as the diversity of leisure activity locations. Spatial patterns show that most individuals—especially active transport users—choose destinations aligned with their SPA-defined opportunity sets, underscoring the metric’s validity in capturing capability sets. Structural equation modeling reveals that SPA directly fosters leisure diversity but also reduces travel time, which in turn is associated with lower diversity. These findings highlight the value of person-centered, capability-informed accessibility metrics for understanding inequalities in urban mobility and informing transport planning strategies that expand real freedoms to participate in social life across diverse population groups.
Keywords Space–time accessibility $\cdot$ Trip chaining $\cdot$ Third-place activities $\cdot$ Structural equation modeling $\cdot$ Transport equity $\cdot$ Human capability approach $\cdot$ Urban mobility behavior
1 Introduction
Transport accessibility is central to inclusive cities through its influence on individual mobility and activity participation (Allen and Farber, 2020; Luz et al., 2022; Liao et al., 2025a). By shaping how easily people can reach different locations and activities, spatial accessibility defines the scope of places and social environments individuals can access (Pereira et al., 2017), particularly to those third-place activities that foster meaningful social interactions (Oldenburg and Brissett, 1982). When transport opportunities are limited, the risk of social exclusion rises for certain groups (Luz and Portugal, 2022; Gallego Méndez et al., 2023). Studies have shown that greater proximity to certain destinations is associated with higher participation rates in those activities (e.g., Althoff et al., 2025). Yet, it remains unclear how transport accessibility, when accounting for individual constraints, links to participation in third-place activities and how this varies across population groups.
Visits to third places shape individuals’ social exposure by linking them to wider society beyond institutional settings, with the extent of such exposure often conditioned by people’s flexibility in allocating time to discretionary activities (Oldenburg and Brissett, 1982). Focusing on leisure activities, as an important part of third-place activities, draws attention to how transport systems must accommodate not only traditional trip purposes, like fixed home-work commuting trips, but also the diverse and time-sensitive mobility demands associated with rich and diverse activities that improve individuals’ well-being (Lee, 2022).
Travel times and time budgets play a central role in shaping individuals’ ability to engage in out-of-home daily activities. A longer travel time enables individuals to extend beyond proximate locations and diversify their activity choices, albeit at the cost of having less available time to engage in activities, whereas tighter time budgets restrict them to a narrower set of opportunities, often closer to home or work anchors. Activity locations in proximity reduce the required travel effort for reaching comparable opportunities, thereby lowering the effective time budget needed to satisfy activity demand. Structural constraints—the layout of the urban amenities and transportation infrastructure—can limit people’s ability to access resources outside the home, hindering efficient time use and participation in broad third-place activities (Luz and Portugal, 2022).
Such structural constraints from the built environment can be largely captured with spatial accessibility measures through network-based models of potential access (Levinson and King, 2020). Nonetheless, most studies in the literature mobilize such transport accessibility measures using home-based and static measures (Ryan et al., 2023), which largely overlook trip chaining and trips that start from non-home destinations, particularly work. Many leisure activities, central to urban vibrancy and the fostering of social interactions (Botta and Gutiérrez-Roig, 2021), take place after work rather than from home (McGuckin and Murakami, 1999). Unlike home- or work-based measures, space-time accessibility metrics account for the spatio-temporal constraints of an individual’s time–space prism (Kim and Kwan, 2003; Victoriano et al., 2020; Saraiva and Barros, 2022), offering a more realistic proxy for their ability to access leisure opportunities.
To date, our understanding of how transport accessibility shapes leisure activity participation remains largely static. Constrained space–time accessibility can reduce mobility efficiency and temporal flexibility, and thereby contribute to limited leisure activity participation. In this study, we i) quantify individual space–time accessibility to leisure opportunities, ii) examine its explanatory power for total travel time and leisure activity participation (location diversity), and iii) assess how these relationships vary across population groups.
1.1 Related work
Third-place activities—such as gym, café, and park—significantly shape social interactions, identity, and community connectedness (Oldenburg and Brissett, 1982). For example, some third places, such as shopping centres, strategically located to connect diverse neighbourhoods, could attract visitors across socioeconomic groups and foster inter-group exposure (Nilforoshan et al., 2023). Institutional locations, i.e., homes and workplaces, serve as anchors that structure daily mobility patterns, constraining the accessibility and relevance of third places (Miller, 1991). The effect of these anchors on mobility is intertwined with individual attributes, lifestyle, and activity demand (Luz and Portugal, 2022).
The proposition that accessibility should be understood as a human capability (Pereira et al., 2017; Luz and Portugal, 2022) offers a unifying perspective that links the multiple factors shaping one’s potential access and observed mobility patterns. This framing consists of two interacting components, the spatial environment and resources, and the individual’s conversion factors. Spatial environment and resources include land-use patterns, transport systems, and their temporal constraints (Liu et al., 2023), while the conversion factors reflect individuals’ perceptions and abilities to translate these resources into real access, i.e., actual travel behavior and activity participation. The conversion factor is shaped by their travel time budget (Chung and Nam, 2024) and cognitive limitations, as well as the wider social, economic, and political environment (Ryan et al., 2019). The interaction of these elements can either expand or restrict people’s capabilities; when restrictions become severe (accessibility poverty), systemic barriers to opportunities lead to transport-related social exclusion (Lucas, 2012), with individuals experiencing different forms of exclusion depending on how spatial resources, personal conversion factors, and external environments combine.
A growing body of evidence shows a strong link between accessibility and realized mobility. Higher accessibility levels are strongly correlated with participation in total, mandatory, and discretionary activities (Fransen et al., 2018; Luz et al., 2022). Better job accessibility is associated with greater mobility ranges (Liao et al., 2025b), and proximity to walkable areas promotes higher levels of physical activity (Althoff et al., 2025). Improving accessibility in neighborhoods with concentrations of low-income or carless households located outside major transit corridors has also been shown to increase daily activity participation (Allen and Farber, 2020). Similarly, enhancing public transit accessibility can particularly benefit discretionary activity participation among disadvantaged groups (Zhang et al., 2022).
Nevertheless, most existing studies rely on home-based cumulative opportunity (Allen and Farber, 2020; Luz et al., 2022) or simple proximity measures (McCormack and Shiell, 2011) of accessibility and treat individual attributes separately. This approach is essentially static and overlooks the mechanisms through which accessibility shapes mobility. For example, while greater job accessibility may be associated with higher leisure participation at the aggregate level (Luz et al., 2022), such a misalignment between the activity types used to measure accessibility and those used to measure participation obscures policy relevance and undermines individual-level insights.
In contrast, space–time accessibility is an individual-based approach that enables the examination of intra-group and intra-location heterogeneities, thereby offering a deeper understanding of accessibility inequalities (Neutens et al., 2010; Saraiva and Barros, 2022). It builds on two core concepts: an individual’s trajectory in space and time (space-time path), and the set of potential trajectories an individual could take given temporal and spatial constraints (anchors) (Kwan, 1998). They define the geographical extent of accessible opportunities (space-time prisms). Measures of space–time accessibility range from direct depictions of prisms (Miller, 1991) to empirically scalable methods based on travel time budgets (Saraiva and Barros, 2022).
Transport accessibility directly shapes travel behavior, with consequences for mobility efficiency—an intermediate factor influencing activity participation. A key expression of efficiency is trip chaining, where individuals combine multiple activities into a single journey rather than returning home between stops (Zhu and Guo, 2022). Recent advances in space–time accessibility incorporate this behavior implicitly by accounting for the feasibility of chaining activities between home and work, given an individual’s travel time budget, residential and workplace locations, and the built environment (Saraiva and Barros, 2022). This formulation reflects realistic mobility patterns, such as engaging in activities on the way home from work, and thus provides a useful proxy for assessing both mobility efficiency and opportunities for third-place participation.
Structural Equation Modeling (SEM) has been widely used in travel behavior research, demonstrating its utility in analyzing complex relationships between the built environment and mobility (Golob, 2003; Kroesen and Van Wee, 2022; Song et al., 2016). SEM is particularly adept at dissecting intricate causal structures (Golob, 2003; Kroesen and Van Wee, 2022). For instance, SEM has been applied to explore the interdependencies among land use, socio-demographic attributes, and travel patterns (Song et al., 2016), and to model mediating variables such as car ownership when analyzing the built environment’s influence on travel behavior (Zhang et al., 2025). This method provides valuable insights into how built environment characteristics—including density, diversity, connectivity, and accessibility—affect various outcomes, such as individual transport emissions, health outcomes, and rural income, by shaping residents’ travel choices and capabilities (Azmoodeh et al., 2023; Kroesen and Van Wee, 2022; Li et al., 2024; Song et al., 2016).
2 Conceptual framework
In this study, we conceptualize space–time accessibility (SPA) as a human capability (Luz and Portugal, 2022), where individuals’ time budgets and mobility resources interact to shape the opportunity sets that on can reach, including third-place activities such as leisure (Figure 1). Conversion factors as a reinforcing cycle: the ability to convert spatial and temporal resources into actual activity participation (functionings) contributes to well-being, which in turn strengthens individuals’ capacity to utilize resources and expand future capabilities.
For example, a person with greater flexibility to integrate social or leisure stops into daily travel (trip chaining) can increase their well-being through richer social interaction, which may enhance confidence and mobility skills that improve their ability to access other opportunities. Conversely, limited time budgets can reduce third-place participation, narrowing social exposure and reinforcing inequalities. Transport planning approaches that equate accessibility solely with the provision of infrastructure or services often overlook these individual-level conversion processes, which are shaped by personal characteristics, environmental barriers, and cultural norms (Pereira et al., 2017). This capability-oriented framework thus shifts attention to how varying levels of space–time accessibility are necessary to ensure equal opportunities, social inclusion, and the freedom to engage in third activities essential for well-being and development.
In this study, we hypothesize that the factors shaping an individual’s participation in leisure activities are those indicated by the thick green arrows (Figure 1). The list of all opportunities accessible by a person (their capability set) is conceptualized as arising from two building blocks: individuals’ home and work/study locations, and their transport and amenity resources.
Figure 1: Accessibility as a human capability. Adapted from (Luz and Portugal, 2022; Luz et al., 2022).
3 Materials
3.1 Study area: The Paris region
The Paris region, commonly known as the Île-de-France, is the most populous of France’s eighteen regions, with an estimated 12.27 million residents as of January 2023 (INSEE, 2022). Centered on the capital Paris in the north-central part of the country, it covers 12,012 km2, about 2% of metropolitan France, yet accounts for nearly 20% of the national population. In the Paris region, population density reaches about 20,200 inhabitants/km2 in the city of Paris, decreases to around 7,000 inhabitants/km2 in the inner suburbs, and averages roughly 3,700 inhabitants/km2 across the wider conurbation (La Grande Conversation, 2023).
The Paris region (see Figure 2), with its monocentric urbanization pattern, has one of the world’s densest and most multifaceted transport networks—spanning multiple transit modes and innovative services (metro, RER, tramways, buses, and on-demand services), coordinated by Île‑de‑France Mobilités (IDFM). The region also offers robust ecosystem support via Mobility-as-a-Service (MaaS) platforms, integrating public transit with shared mobility, MaaS apps, and seamless ticketing (Île-de-France Mobilités, 2023). Residents feature frequent multimodal trips and rich variation across individuals (Yin and Leurent, 2023).
Figure 2: The Paris region. Black dots in the bottom-right corner of the figure indicate anchor locations within the highlighted box.
3.2 Trip records
We use a dataset collected via the Enquête Mobilité par GPS (EMG 2023) initiative and made available through participating in the NetMob 2025 Data Challenge (Chasse et al., 2025). The EMG 2023 survey, conducted between October 2022 and May 2023 covered 3,337 residents aged 16–80 in the Paris region, excluded non-residents, tourists, and the immobile, and used a multichannel recruitment strategy combining quotas and random draws to ensure representativeness (Chasse et al., 2025).
In this study, we analyze 14,169 individual-day records from 2,415 individuals with identified home and work locations, including information on date, day type (e.g., strike, holiday), origin, destination, transport mode, trip duration, and purpose. Origins and destinations are represented by the centroids of the visited locations’ corresponding H3 hexagons at resolution 10 (Uber Technologies, Inc., 2018), each covering approximately 0.015 km2.
3.3 Socio-economic attributes
The Individuals dataset provides a rich set of sociodemographic, household, and mobility-related attributes. It includes basic identifiers and residence information such as municipality codes, as well as demographic variables like sex, age, education level, and socio-professional category. We also augment the attributes with the poverty rate at the IRIS (9-digit) zone level, used as a proxy for income in 2021. This measure corresponds to the share of individuals living in households whose standard of living—after accounting for taxes, transfers, and social benefits—falls below 60% of the national median disposable income (Institut national de la statistique et des études économiques (2024), INSEE). Household composition is captured through the number of persons and their age distribution, along with housing type. Mobility resources and constraints are represented by indicators of driving licence ownership and car availability, complemented by information on access to two-wheelers, bicycles, e-scooters, and other mobility devices. The dataset also records subscriptions to public transport and other mobility services, alongside a statistical weighting coefficient for representativeness.
We focus on the main attributes: age, gender, poverty rate (zone level), education, household structure, main transport mode, active mode use, and public transport subscription.
3.4 Public transit and road network data
For public transit, we collected GTFS schedules from transport.data.gouv.fr, the French national open data platform for mobility, which provides official timetables and service information across transit operators (IDFM, accessed on 30 June 2025). Road network data were obtained from OpenStreetMap, an open, collaboratively maintained geospatial database offering detailed and regularly updated representations of streets, pathways, and transport infrastructure (accessed on 1 July 2025).
3.5 Point of interest (POI) data
POI data were collected via the Overture API, restricted to the geographical extent of the trip records and filtered for a confidence level above 0.7 (Overture Maps Foundation, 2023). We then labeled these POIs into two categories aligned with the activity purposes recorded in the trip data, Social & Leisure (67 k), consolidated through a combination of GPT-4o–assisted classification and manual validation.
4 Methodology
4.1 Space-time accessibility
In this study, we operationalize individual-based capability set as space–time accessibility (SPA). Specifically, we focus on the cardinal accessibility dimension, which accounts for the number of opportunities (leisure POIs, including parks, etc.) that individuals can feasibly reach within their available time budget (Saraiva and Barros, 2022).
Figure 3: Space–time accessibility (SPA). Adapted from Saraiva and Barros (2022).
The measure builds on the space-time feasibility concept (Kwan, 1999) and is modified based on the cardinal individual-based accessibility measure proposed in (Saraiva and Barros, 2022). Each individual is assumed to have a fixed time budget $tb_{i}$, defined as the residual time available for discretionary activities after completing mandatory activities (e.g., at home and work). An activity $k$ is considered accessible to individual $i$ as if he/she does a one-stop trip chaining in between their trip from work to home, as more commonly found in the population (McGuckin and Murakami, 1999). In other words, the total time required to (i) travel from home to work ($t_{hw}$), (ii) travel from work to $k$ ($t_{wk}$), and (iii) return from $k$ to home ($t_{kh}$) does not exceed the time budget.
$a^{k}_{i}=\begin{cases}1,&\text{if }t_{hw}+t_{wk}+t_{kh}\leq tb_{i}\\ 0,&\text{otherwise}\end{cases}$ (1)
where:
  • $t_{hw}$ is the travel time from home to work,
  • $t_{wk}$ is the travel time from work to activity location $k$,
  • $t_{kh}$ is the travel time from activity location $k$ to home,
  • $tb_{i}$ is the time budget of individual $i$.
The overall space–time accessibility $A_{i}^{k}$ of individual $i$ is then given by the sum of all feasible opportunities under the $k$ category:
$A_{i}^{k}=\sum a^{k}_{i}$ (2)
In this study, we define the individual travel time budget $tb_{i}$ as 90 minutes for all individuals. Commuting time ($t_{hw}$) is empirically estimated using the weighted median of observed commuting trips. Discretionary activity $k$ is defined as a visit to leisure-related locations (e.g., bars).
For calculating $t_{wk}$ and $t_{kh}$, we fix the departure time at 17:00, resembling the work-to-home trip chaining. Travel times between home, work, and candidate leisure POIs were computed for both public transit and car, using GTFS schedules and road network data, implemented through the r5r package (Pereira et al., 2021).
4.2 Capability set vs. spatial choices
While the space-time accessibility construct remains somewhat opaque, realistic in formulation, the actual spatial choices in leisure activity participation can differ. To gain a finer-grained understanding of how accessibility structures behavior, we assess whether individuals tend to select activity locations that align with what their modeled opportunity landscape (SPA) predicts as desirable, or whether actual behavior departs from this logic.
This behavioral analysis focuses on the alignment between the ranked set of feasible locations SPAi (adding up to $A_{i}$) and the actual visited locations for each individual $i$. By comparing observed choices against random draws from SPAi, we test for selectivity—that is, whether individuals systematically prefer higher-ranked locations.
We implement a rank-based comparison. We evaluate whether individuals tend to choose better-ranked locations (e.g., those that take a shorter time to reach) than would be expected by chance. For each individual $i$, we construct their SPAi—a ranked list of feasible locations (H3 hexagons at resolution 8, $\sim$0.74 km2) sorted from 1 (best) to $N_{i}$ (worst) based on the modeled opportunity score (e.g., time left in the travel time budget $tb_{i}-t_{hw}-t_{wk}-t_{kh}$).
We then compare the actual visited locations $V_{i}$ to this ranking. First, we log the share of locations in $V_{i}$ that fall outside SPAii. Second, the test statistic is the average rank of these visited locations fall within SPAii:
$T_{\text{mean},i}^{\text{act}}=\frac{1}{k_{i}}\sum_{\ell\in V_{i}}\text{rank}(\ell)$ (3)
To generate a null distribution, we perform $B=1000$ random draws of $k_{i}$ locations from SPAii, computing $T_{\text{mean},i}^{(b)}$ for each. We then derive both the empirical $p$-value:
$p_{i}=\frac{1+\#\left\{b:T_{\text{mean},i}^{(b)}\leq T_{\text{mean},i}^{\text{act}}\right\}}{B+1}$ (4)
and the standardized effect size:
$d_{i}=\frac{T_{\text{mean},i}^{\text{act}}-\mu_{i}^{\text{null}}}{\sigma_{i}^{\text{null}}}$ (5)
where $\mu_{i}^{\text{null}}$ and $\sigma_{i}^{\text{null}}$ are the mean and standard deviation of the null distribution. Negative values of $d_{i}$ indicate better-than-random performance—i.e., that the individual tends to visit higher-ranked locations than would be expected under random choice.
4.3 Mobility quantification
We quantify individual mobility along two dimensions: trip making, captured by daily total travel time, and third-place activity participation, characterised by the location diversity. Because the travel diary lacks detailed labeling of third-place activities, we approximate third-place participation using all leisure activities (’LEISURE’). We compute a diversity measure based on the distribution of leisure activity locations. For each individual $i$, we count the number of visits to each distinct location. Let $n$ denote the total number of leisure visits, and let $K$ denote the number of distinct locations visited. Defining $p_{k}=n_{k}/n$ as the relative frequency of visits to location $k$ ($k=1,\dots,K$), we compute the Hill number of order $q=1$, which is given by:
$H_{1i}=\exp\left(-\sum_{k=1}^{K}p_{k}\,\ln(p_{k})\right).$ (6)
which reflects the effective number of equally frequent locations visited, and accounts for both richness and evenness in the individual’s leisure activity location distribution.
4.4 Structural equation modeling specification
Based on the data summarized in Table 1, we estimate a structural equation model (SEM) linking (i) individual attributes, including one lifestyle metric on active transport mode use, (ii) transport mode, (iii) space-time accessibility (SPA), (iv) trip making, and (v) leisure activity participation. Let $i$ index individuals. The observed variables are:
  • $A_{i}$: log–transformed space–time accessibility to leisure opportunities;
  • $\mathbf{Z}_{i}$: a vector of exogenous sociodemographic and lifestyle attributes (dummy–coded for household type, gender, education, etc.), retained after preprocessing to remove colinearity and low variance;
  • $\mathbf{M}_{i}$: transport variables, including the main transport mode dummy and the public transport subscription dummy;
  • $B_{i}$: travel behavior indicator, namely total travel time;
  • $H_{1i}$: leisure activity participation, measured by the Hill number of order $q=1$.
Table 1: Descriptive statistics of the 2,415 commuting individuals. Values represent weighted means (continuous variables) and weighted percentages (categorical variables).
Group Variable Levels Mean (SD) or %
Individual attributes Age - 43.3 (11.8)
Poverty rate (IRIS zone level) - 15.8 (8.3)
Education - No diploma 1.7
- Vocational 11.7
- Lower secondary 3.0
- Upper secondary 31.4
- 3–4-year higher education 15.7
- 5-year-and-above higher education 25.6
- Missing 10.9
Gender - Man 47.6
- Woman 52.4
Household type - Living alone 12.2
- In a couple w/o children 20.3
- Single parent 11.6
- Living with parent(s) 6.2
- Not related to other household members 1.1
- In a shared apartment 0.3
- In a couple w/ child(ren) 47.2
- Another family member in the household 1.1
Use of active mode - No 62.1
- Yes 37.9
Space-time accessibility Space-time accessibility (>0) - Non-zero 39.7
- Zero 60.3
Space-time accessibility (log) - 2.9 (3.8)
Transport mode Main transport mode - Car 38.7
- Public transit 61.3
Public transit subscription - No 35.7
- Yes 64.3
Trip making Total travel time (min) - 96.2 (44.1)
Leisure activity participation Leisure activity diversity - 14.6 (11.9)
Before estimating the structural equation model, it is important to specify a directed acyclic graph (DAG) that encodes the theoretical and causal assumptions about the relationships among the variables. The DAG serves as a transparent foundation for identifying potential confounding paths, testing conditional independence implications, and ensuring that the SEM structure is both theoretically grounded and empirically justified.
After testing for testable implications, the final directed acyclic graph (DAG) reflects the theoretically informed and empirically tested structure of the relationships among individual attributes, transport mode, capability set, trip making, and activity participation (Figure 4). Individual characteristics are modeled as exogenous factors that shape both transport mode choice and accessibility (capability set), as well as exerting direct influences on trip-making and activity participation. Transport mode mediates part of these effects by influencing accessibility, trip-making, and participation directly. The capability set, representing the feasible space–time leisure opportunities available to individuals, is affected by both personal attributes and transport mode, and in turn drives patterns of trip-making and participation. Trip-making operates as an intermediate behavioral mechanism linking accessibility to leisure activity participation. Finally, leisure activity participation is conceptualized as the outcome of multiple interacting pathways: directly from individuals, from transport behavior, and indirectly via both the capability set and trip-making. This structure balances theoretical expectations from time-geography and accessibility theory with empirical evidence from conditional independence testing, ensuring that only statistically supported pathways are retained.
Individualattributes($\mathbf{Z}_{i}$)Transportmode($\mathbf{M}_{i}$)Capability Set($A_{i}$)Trip making($B_{i}$)Activityparticipation($H_{1i}$)
Figure 4: Pathway structure. This directed acyclic graph shows the hypothesized and empirically validated pathways. Exposure=Capability set, Outcome=Activity participation, Black arrows=Causal paths, Brown arrows=Biasing paths.
In SEM, the structural relations in Figure 4 are specified as follows:
$M_{i1}$ $\sim\mathbf{Z}_{i},$ (7)
$M_{i2}$ $\sim\mathbf{Z}_{i},$ (8)
$A_{i}$ $\sim\boldsymbol{\alpha}_{Z}^{\top}\mathbf{Z}_{i}+\boldsymbol{\alpha}_{M}^{\top}\mathbf{M}_{i},$ (9)
$B_{i}$ $\sim a_{tt}\,A_{i}+\boldsymbol{\gamma}_{Z}^{\top}\mathbf{Z}_{i}+\boldsymbol{\gamma}_{M}^{\top}\mathbf{M}_{i},$ (10)
$H_{1i}$ $\sim c\,A_{i}+b_{tt}\,B_{i}+\boldsymbol{r}_{M}^{\top}\mathbf{M}_{i}+\boldsymbol{d}^{\top}\mathbf{Z}_{i}.$ (11)
The model captures both direct associations of SPA with leisure participation ($c$), and indirect associations transmitted through travel behavior pathways ($a_{tt}b_{tt}$). In addition, sociodemographic attributes $\mathbf{Z}_{i}$ contribute to leisure activity location diversity both directly (coefficients $\mathbf{d}$) and indirectly via transport mode ($\mathbf{M}_{i}$), SPA ($A_{i}$), and trip making ($B_{i}$).
The analysis was performed in R (lavaan 0.6-19) using the diagonally weighted least squares (DWLS) estimator, which is appropriate for models that include categorical or non-normally distributed variables. Continuous variables were standardized to stabilize variances. Weights were normalized to have a mean of one and applied during estimation to account for individual-level representativeness. To detect multicollinearity, the variance inflation factor (VIF) analysis was applied to evaluate all the candidate variables. We removed the variable Age with a VIF above 7. Near-constant and collinear dummy variables were further pruned to avoid estimation instability (Education, Gender, and Poverty rate). We report standardized coefficients, $R^{2}$ values for endogenous variables, and standard SEM fit indices (CFI, TLI, RMSEA, SRMR). Indirect, direct, and total associations are computed through model-implied parameter combinations.
5 Results and discussion
5.1 Descriptive statistics
Figure 5 presents a spatial overview of space–time accessibility (SPA), trip making, and leisure activity participation across the study region, highlighting the geographic variations. Figure 5a shows the proportion of inhabitants with SPA >0>0, indicating the share of individuals who have feasible opportunities beyond their fixed home and work locations. Higher values are concentrated in the northern central urban areas and scattered across select suburban zones, suggesting greater access to leisure activity opportunities in both dense and well-connected peripheral locations, mainly by car. Figure 5b maps the average total travel time, with higher values concentrated in peripheral areas. This pattern underscores the role of trip-making as a necessity in facilitating activity demand rather than a choice, driven by limited service availability on the outskirts. Figure 5c visualizes leisure activity location diversity. Compared to SPA and total travel time, diversity values show weaker spatial clustering, as they are more strongly shaped by individual-level differences in activity behavior. Together, the maps suggest that while core urban areas tend to offer greater capability sets and diverse leisure opportunities, spatial disparities in accessibility, mobility, and activity participation persist across individuals.
Figure 5: Space–time accessibility (SPA) and mobility metrics on the map. a, Share of inhabitants with SPA > 0. b, Total travel time (daily, min). c, Leisure activity diversity (weighted median of each zone’s inhabitants).
We further examine how SPA relates to total travel time and leisure activity participation, stratified by main transport mode (car vs. public transit) in Figure 6. Figure 6a shows that individuals with SPA $\neq 0$ have substantially lower travel times, suggesting that feasible opportunities beyond home and work enable more efficient activity engagement. Figure 6b illustrates a clear negative association between SPA and travel time for both car and public transit users, with the effect more pronounced among car users. Figure 6c shows that leisure activity diversity increases with accessibility, particularly for public transit users. Taken together, these results indicate that higher accessibility not only facilitates efficiency but also broadens the diversity of leisure participation, though the magnitude and pathways of these effects differ between car and public transit users.
The correlational results (Figure 6) support our assumptions regarding the interdependence of individual attributes, transport mode choice, and space–time accessibility (Figure 2). They further suggest that an individual’s space–time accessibility shapes both trip-making behavior and actual leisure activity participation. SEM outcomes further disentangle their relationships.
Figure 6: Space–time accessibility $A_{i}$ and trip making and activity behaviors across transport modes. a, Total travel time and leisure activity diversity, comparing those with $A_{i}$ = 0 (grey) and $A_{i}$ > 0 (green). b, Total travel time as a function of $A_{i}$ for car and public transit users ($A_{i}$ > 0). c, Leisure activity diversity ($A_{i}$ > 0) by mode. Error bars indicate bootstrap median estimation errors.
5.2 Selectivity in spatial choice of leisure activities
Based on 856 individuals where their $A_{i}>0$ and have leisure activities observed, we show a strong alignment between SPAi and actual choices of leisure activities (Figure 7). Only 0.1% of individuals have zero visited places within their capability set, the distribution of visited places outside SPAi is skewed toward lower values (Figure 7a): with a bottom 10% who are mainly car users having over 50% of their visits outside their capability set. Individuals who frequently travel beyond their SPA are more likely to rely on cars. Space-time accessibility in this context primarily captures trip-chaining opportunities between home and work, framing leisure visits as secondary or incidental. While informative, this representation can be inadequate for modeling car users who travel for longer than 90 minutes, whose greater spatial flexibility allows for more diverse and decoupled travel patterns.
Most individuals exhibit negative $d_{i}$ values (Figure 7b), indicating a tendency to choose better-ranked locations than expected by chance, as 67% of individuals have significant deviation from being random ($p<0.05$). This supports the feasibility of using SPA as a realistic approximation of individuals’ capability sets. A cluster of individuals shows strongly negative $d_{i}$ with highly significant $p$-values, while others have values near zero or even positive, suggesting closer-to-random or worse-than-random behavior. These results reveal considerable heterogeneity in the degree of behavioral selectivity across the population.
Among all individual attributes considered, only the use of active transport mode shows a statistically significant association with the effect size ($d_{i}$), as illustrated in Figure 7c. Non-active users exhibit a higher median $d_{i}$ and a tighter distribution skewed toward weaker selectivity (i.e., more positive values). In contrast, active mode users show stronger selectivity as expected by their capability set ($\beta=-0.51,p<0.001$). These results suggest that individuals using active mode tend to align more closely with their SPA-based opportunity rankings, whereas car users may be less constrained or freer in their spatial choices.
Individuals exhibit varying degrees of alignment with modeled opportunity structures. Car users have a higher chance of traveling beyond their capabilities. Meanwhile, active users show stronger within-SPA selectivity. These findings underscore the value of decomposing behavioral mechanisms underlying accessibility: individuals do not uniformly act upon modeled opportunities, and their selectivity depends critically on their transport mode and spatial constraints.
Figure 7: Selectivity in spatial choice behavior across individuals. a, Distribution of individuals by the share of visited locations outside their modeled feasible set (SPAi), with bar color indicating the proportion of car users in each bin. b, Standardized effect size $d_{i}$ plotted against the empirical $p$-value from the mean rank test, on a log scale. The vertical line marks the $p=0.05$ threshold. Each point is an individual. c, Boxplot of $d_{i}$ stratified by use of active transport mode, with jittered points colored by statistical significance (blue: $p<0.05$).
5.3 Modeling outcomes
We apply structural equation modeling techniques with the diagonally weighted least squares (DWLS) estimator on a total of 2,415 weighted observations (individuals). The model demonstrated excellent overall fit, with a Comparative Fit Index (CFI) of 0.998 and Tucker–Lewis Index (TLI) of 0.963. The Root Mean Square Error of Approximation (RMSEA) was 0.041, with a 90% confidence interval of [0.018, 0.067], and the Standardized Root Mean Square Residual (SRMR) was 0.008, all supporting model adequacy (Hu and Bentler, 1999).
Significant direct and indirect pathways with standardized coefficients exceeding an absolute value of 0.1 (Figure 8) highlight the role of household structure, active mode use, and public transport subscription in shaping space–time accessibility, trip making, and leisure activity outcomes.
Figure 8: Structural equation model of factors shaping leisure activity participation. Standardized path coefficients ($p<0.05$) are shown along the arrows, with positive effects in teal and negative effects in orange. Individual attributes (left) influence transport mode choices and space–time accessibility (center), which in turn shape travel behavior (trip making) and leisure activity participation (right).
5.3.1 Transport mode and space-time accessibility
Household structure and lifestyle indicators emerged as strong predictors of transport mode choice. In particular, living in a couple with children substantially increases the likelihood of using the car as the main mode ($\beta=0.17$, $p<.001$), which aligns with prior evidence that households with children face greater time constraints and logistical needs, thereby relying more heavily on private vehicles (Scheiner and Holz-Rau, 2013). In contrast, individuals reporting active mode use are much less likely to depend on the car ($\beta=-0.63$, $p<.001$).
Public transport subscription also shows important interactions with lifestyle and household type. While subscription is positively associated with active mode use ($\beta=0.11$, $p<.001$), reflecting complementarities between walking/cycling and transit (Oeschger et al., 2020), it is negatively associated with family-related household types, such as couples with children ($\beta=-0.19$, $p<.001$). This suggests that households with children are less able to rely on public transit, possibly due to constraints of escorting children and multi-stop trip chaining (Scheiner and Holz-Rau, 2013).
Space-time accessibility is shaped by both sociodemographic and transport characteristics. Active mode users—those regularly walking or cycling—exhibit significantly higher SPA ($\beta=0.17$, $p<.001$). Given how SPA is evaluated, this could be ascribed to two main mechanisms: (1) these individuals mostly live in the city of Paris, with good proximity to leisure opportunities and good support for walking and cycling; and (2) they have relatively short commuting times (34 min) compared to non-active mode user (43 min), making the assumed 90-minute daily travel time budget have a greater share available for potential leisure activities.
The positive relationship between car usage and STA ($\beta=0.29$, $p<.001$), while public transport subscription is negatively associated with STA ($\beta=-0.18$, $p<.001$). This suggests that, given individuals’ fixed daily anchors (home and work/study locations) and their chosen mode of travel, the car provides more effective access to a wider range of leisure opportunities compared to public transport.
Household structure itself significantly shapes accessibility. Respondents living in couples with children yield lower STA ($\beta=-0.1$, $p=.001$). One explanation is that caregiving responsibilities often reduce spatial and temporal flexibility, leading to a tight travel time budget for planned activities.
5.3.2 Trip making
Space-time accessibility strongly predicts total travel time ($\beta=-0.37$, $p<.001$). Good SPA implies a greater number of accessible leisure opportunities within the time–space prism, which in turn leads to shorter travel times, indicating that part of the activity demand can be met in proximity. The use of active modes is also associated with shorter total travel time, suggesting that those who rely on active transport do so because their activity needs can be met closer to their home and work anchors. The total travel time, in turn, serves as a mediator between SPA and leisure activity participation.
5.3.3 Leisure activity participation
The leisure activity diversity is shaped by space-time accessibility, travel behaviors, and individual attributes. Higher accessibility is positively associated with greater location diversity ($\beta=0.13$, $p<.001$). This finding underscores a conceptual alignment between space-time accessibility to leisure opportunities and actual participation in diverse leisure activities. In other words, individuals with higher accessibility to leisure opportunities are more likely to translate potential into visits to a broader range of leisure locations.
Trip making also contributes here: longer travel times increase leisure activity time ($\beta=0.23$, $p=<.001$). From the perspective of utility-maximization, longer travel time can still result in higher overall utility if it enables meaningful leisure engagement (De Vos et al., 2016). This effect is even stronger than that of SPA directly, underscoring the role of total travel time as a key mediator shaping actual leisure activity participation.
Lifestyle differences are notable. Active mode use is positively associated with leisure activity location diversity ($\beta=0.23$, $p<.001$). Leisure travel is often motivated by social needs, the pursuit of variety, and personal recharge (Stauffacher et al., 2005). Active transport may facilitate these motivations by providing greater flexibility in route selection and opening up opportunities to access varied leisure spaces such as parks, cafés, or community centers. It’s plausible that walking or cycling itself becomes part of a user’s leisure experience—adding intentionality and value to the travel act.
5.3.4 Effect decompositions of space-time accessibility
Decomposition analyses show that space–time accessibility had both direct and indirect effects on leisure participation (Table 2). The direct effect of accessibility on leisure diversity is positive and significant ($\beta=0.13$, $p<.001$), while the indirect effect through travel time is negative and significant ($\beta=-0.09$, $p<.001$). As a result, the total effect is small and marginally statistically significant ($\beta=0.04$, $p=0.077$), indicating that positive direct influences are offset by negative indirect pathways. In addition, active mode use exerted strong positive direct effects on leisure participation ($\beta=0.24$, $p<.001$), underscoring heterogeneity in activity outcomes beyond accessibility constraints.
Table 2: Estimation results of SEM (significant at $\alpha=0.05$). Standardized coefficients reported.
Direct Indirect Total
Effect on total travel time
   SPA ($A_{i}$) $^{***} 0.37-0.37^{***}$
   Active mode $^{***} 0.12-0.12^{***}$
Effect on leisure diversity
   SPA ($A_{i}$) $13^{***} 0.0$$8^{***} 0.0$40.04
   Travel time $^{***} 0.230.23^{***}$
   Active mode $^{***} 0.240.24^{***}$
   Car use (main mode) $^{**} 0.180.18^{**}$
   PT subscription $0.035$ $0.035$
Notes: SPA = space–time accessibility. Indirect effect of SPA operates through total travel time. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
While the total effect of SPA on leisure participation was not strongly significant ($p=0.077$), the decomposition highlights that SPA does matter: it exerts a positive direct effect on activity diversity, but this is offset by a negative indirect effect operating through reduced travel time. These countervailing mechanisms suggest that accessibility shapes leisure participation in complex ways, rather than through a simple net effect.
6 Conclusion
This study investigated how individual-based space–time accessibility (SPA) influences leisure activity participation, with a focus on the Paris region. By leveraging high-resolution mobility data, multimodal transport networks, and a capability-based accessibility framework, we demonstrated that SPA—operationalized as the feasible set of leisure opportunities within individuals’ time budgets—plays a critical role in shaping both the efficiency of travel (via total travel time) and the diversity of leisure activity locations.
Our findings provide strong empirical support for the hypothesis that higher SPA enables more efficient traveling in meeting daily activity demand and richer leisure activity engagement. Through structural equation modeling, we identified both direct and indirect pathways: SPA directly promotes greater leisure diversity, but it also reduces travel time, which is associated with lower diversity, resulting in offsetting effects. The observed behavioral selectivity in spatial choices further validated the SPA construct as a realistic proxy of individuals’ capability sets—particularly among active transport users, who showed stronger alignment with modeled opportunity structures.
Importantly, the study revealed substantial heterogeneity in accessibility and activity outcomes across population groups. Household structures involving caregiving responsibilities (e.g., single parents, couples with children) display lower SPA and reduced leisure participation, pointing to persistent structural constraints. These findings underscore the importance of considering both the spatial distribution of opportunities and the temporal feasibility of access when designing inclusive transport policies.
By embedding accessibility into a human capabilities framework, this work shifts the discourse from infrastructure provision toward understanding how individuals convert transport resources into meaningful participation. Future research should extend this analysis to non-commuting populations, incorporate dynamic time budgets, and examine the longitudinal impacts of accessibility improvements on social inclusion and well-being. As cities pursue more equitable and sustainable mobility systems, metrics like space-time accessibility can serve as valuable tools to identify gaps, monitor policy impacts, and prioritize interventions that expand real freedoms to engage in everyday urban life.
Data availability
The data used in this study were provided through participation in the NetMob 2025 Data Challenge (Chasse et al., 2025), under a non-disclosure agreement (NDA) between all authors, Inria with IFPEN. Due to licensing terms and privacy constraints governed by the European General Data Protection Regulation (GDPR), access to the data is restricted. Venue locations and categories can be retrieved from Overture API. Census data (income) were collected from Institut national de la statistique et des études économiques (INSEE) that is publicly available. GTFS data were collected from transport.data.gouv.fr that are publicly available. All data were utilized in accordance with the terms of service specified by their respective provider.
We adhered to the guidelines by the Chalmers Institutional Review Board (IRB) according to the Swedish Act (2003:460) concerning the ethical review of research involving humans, as well as the General Data Protection Regulation 2016/679 (GDPR). According to the data applied, the study was exempt from ethical review under the Swedish Ethical Review Act (2003:460).
Python (version 3.11) code and R (version 4.5.1) code were used to analyse and visualize the data. The accessibility-related travel times were calculated using r5r (version 2.3.0). Code to reproduce our results is publicly available on GitHub Repository.
Acknowledgements
This research is funded by the Swedish Research Council (Project Number 2022-06215).
Author contributions
Y.L. conceptualized the study. All authors designed the methods. Y.L. processed the data and the model. All authors wrote the manuscript.
Competing interests
The authors declare that there are no conflicts of interest.
Additional information
Correspondence and requests for materials should be addressed to Y.L.